OpenStudy (anonymous):

B is the midpoint of. AB=3x-6 BC=2x+2

OpenStudy (anonymous):

are we finding x again

OpenStudy (anonymous):

what is b the midpoint of?

OpenStudy (anonymous):

of AC with the line on top

OpenStudy (anonymous):

What are we finding the value of?

OpenStudy (anonymous):

the value of AB

OpenStudy (anonymous):

okay so you still set it up like this 3x-6 = 2x+2

OpenStudy (anonymous):

find x

OpenStudy (anonymous):

can you do that yourself I have walked you though that a couple of times

OpenStudy (anonymous):

i got 5x= 8

OpenStudy (anonymous):

3x - 2x - 6 = 2x - 2x +2 x - 6 = 2 x - 6 + 6 = 2+ 6 x = 8

OpenStudy (anonymous):

do you see what I did?